# What Do Latent Action Models Actually Learn?

### v1

@@ Abstract
Latent action models (LAMs) aim to learn action-relevant changes from unlabeled videos by compressing changes between frames as latents. However, differences between video frames can be caused by controllable changes as well as exogenous noise, leading to an important concern – do latents capture the changes caused by actions or irrelevant noise? This paper studies this issue analytically, presenting a linear model that encapsulates the essence of LAM learning, while being tractable. This provides several insights, including connections between LAM and principal component analysis (PCA), desiderata of the data-generating policy, and justification of strategies to encourage learning controllable changes using data augmentation, data cleaning, and auxiliary action-prediction. We also provide illustrative results based on numerical simulation, shedding light on the specific structure of observations, actions, and noise in data that influence LAM learning.
@@ 1 Introduction
Latent action models (LAMs) aim to infer controllable action changes from streams of image observations in an unsupervised manner (Rybkin et al., 2018; Menapace et al., 2021). This is valuable because action-labeled data is typically expensive to source, while unlabeled videos are abundant. Hence, it offers a route for embodied AI systems to learn from large unlabeled datasets, for instance using the inferred latent actions as targets for pre-training a policy, while a small amount of labeled data can be used to learn a mapping from latent to real action controls (Ye et al., 2024). This has proven effective in learning from videos of 2D video games, robotics, and even broadcast tennis footage (Menapace et al., 2021; Schmidt and Jiang, 2023; Bruce et al., 2024; Chen et al., 2024b; Ye et al., 2024; Sun et al., 2024; Cui et al., 2025; Gao et al., 2025).
@@
The success of such LAM-based recipes relies on the inferred latent action labels mapping to the real control action signals of interest. However, there is concern that this may not always be the case. For example, there is an intuition that LAM’s inferred latent ‘actions’ simply compress differences between consecutive frames, even when control actions are not the cause of those differences (McCarthy et al., 2024). As such, LAMs may only succeed in domains where the cause of changes between observations can be fully attributed to the control action. Another example is that, although bottleneck is widely used in different LAMs, whether it necessary or whether it can eliminate the dependence of the latent on the observation is unclear (Schmidt and Jiang, 2023).
@@
Figure 1: Linear LAM is an abstraction of the LAMs used in previous work. Inputting consecutive observation pairs $(\boldsymbol{o},\boldsymbol{o}^{\prime})$, the LAMs output the second observation via a reconstruction loss, $\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert_{2}^{2}$. An information bottleneck tries to stop the direct copying of $\boldsymbol{o}^{\prime}$, with the expectation the latent $\boldsymbol{z}$ will correspond to the control action $\mathbf{a}$. Linear LAM captures the essence of LAM training whilst being analytically tractable. The diagrams of previous LAMs are copied from their original papers: LAPO (Schmidt and Jiang, 2023), LAPA (Ye et al., 2024), Moto (Chen et al., 2024c), Genie (Bruce et al., 2024), AdaWorld (Gao et al., 2025), and Go-1 (AgiBot-World, 2025).
@@
To elucidate LAMs, this paper conducts a theoretical analysis of a linear version of LAM. By retaining the architecture of recent LAM models, but swapping deep neural network components with simpler linear layers, we are able to preserve the fundamental challenge of LAM training but in an analytically tractable form. Our analysis of linear LAM provides precise insights into, e.g., when inferred latent actions capture true control signals, and when they don’t, and what information is captured by different components within LAM. Surprisingly, our analysis also reveals additional issues not currently known to the LAM community – related to the over-parametrization property of LAM and the randomness of data-generation policy. Finally, we propose and study potential solutions to ameliorate these issues – data augmentation and predicting action as an auxiliary task.
@@
    (Section 3) We present linear LAM, a framework that captures the essence of LAM in practice while being tractable.
@@
    (Section 4.1) We show linear LAM reduces exactly to principal component analysis (PCA) on a mixture of controllable changes and exogenous noise, under certain assumptions. Our analysis justifies the general LAM framework by showing that LAM is possible to capture the true action without accessing action labels.
@@
    (Section 4.2) Correlations between observations and actions in data reduces LAM’s focus on the controllable changes. This suggests that higher randomness in data-generating policies benefits LAM’s learning.
@@
    (Section 4.3) We validate that performing data augmentation during LAM training can mitigate the over-parametrization issue and thus improve the semantics of the latent.
@@
    (Section 4.4) We find that adding an action-prediction head encourages LAM to prioritize the learning of controllable changes for the latent.
@@ 2 Related Work
A similar topic with longer history is learning action representations in reinforcement learning (RL). For instance, PG-RA (Chandak et al., 2019) clusters actions based on the similarity of their impact on the state to improve generalization in the action space. LASER (Allshire et al., 2021) learns latent action through an encoder-decoder architecture to reconstruct real actions, resulting in higher learning efficiency in RL. TAP (Jiang et al., 2022) learns the latent action that can help to reconstruct the full trajectory (with state, action, and reward) condition on the state. EAR (Hua et al., 2022) finds that the latent task embedding resulting from the training of multi-task policies emerges to be good action representation with a gometrically and semantically meaningful structure. AD3 (Wang et al., 2024) adopts a inverse dynamics model and a forward dynamics model to extract latent action, similar to popular LAMs, but conditions these models on the real actions.
@@
Different from the above papers that learn action representation based on real actions, our paper focuses on the model where the latent action is learned without access to the real action labels. Without the need for action labels during training, it is advantageous as it allows leveraging Internet-scale video datasets (Miech et al., 2019; Chen et al., 2024a; Pei et al., 2025; Wang et al., 2023) in the pre-training stage, which contain rich expert demonstrations but lack action labels. While high-quality robotic datasets with action annotations exist (Vuong et al., 2023; Fang et al., 2023; Khazatsky et al., 2024; AgiBot-World, 2025), they remain limited in scale. Such datasets can be integrated into the semi-supervised learning framework to extract latent actions (Nikulin et al., 2025) or the policy fine-tuning stage (Schmidt and Jiang, 2023; Ye et al., 2024). Although some researchers try to extract pre-defined actions from observations using computer vision techniques (Mendonca et al., 2023), latent action models (LAMs) can extract latent actions to represent more general transformation.
@@
Information bottleneck or auto-encoder is adopted as the core technique in training LAMs in an unsupervised learning manner since Rybkin et al. (2018). ILPO (Edwards et al., 2019) learns latent actions along with the training of a policy that outputs the latent action and a world model that conditions on the latent action. Menapace et al. (2021) proposes a probabilistic action network that extracts a discrete action label and a continuous action variability embedding from consecutive observations. This network is trained jointly with an action decoder to generate video controlled by extracted actions. While these work adopts bottleneck in the learning of latent actions, their training losses are complicated by involving policy learning or recurrent networks. LAPO (Schmidt and Jiang, 2023) proposes that LAM design with a inverse dynamics model that extract the latent action and a forward dynamics model that reconstruct the next observation based on the latent action. Starting from this paper, many papers follow quite similar architectures to LAPO but target for a discrete latent action space such as FICC (Ye et al., 2022), Genie (Bruce et al., 2024), LAPA (Ye et al., 2024), Moto (Chen et al., 2024c), IGOR (Chen et al., 2024b), Go-1 (AgiBot-World, 2025), and GR001T N1 (Nvidia, 2025). However, there is a debate whether discrete latent actions are better than continuous latent actions. For example, Nikulin et al. (2025) finds that using continuous latent actions with larger bottleneck can not only result in better predictability on real actions but also lead to better performance for downstream policies. We consider continuous latent action in our paper and leave the analysis on vector quantized latent action in future work. In contrast to the popularity of LAM in recent foundation models for embodied AI, issues around the objective, learnability, and robustness of LAM have received limited attention. Consequently, our paper aims to elucidate these questions.
@@ 3 Setup
This section broadly introduces the problem setting, goal and model used in recent LAM work in practice. We then more formally detail these for linear model to be used in subsequent analysis. At last, we outline the details of our simulation setup to be used in support of our later analysis.
@@ 3.1 LAMs in Practice
Setting. The setting tackled by recent LAM work assumes access to a large dataset of pairs of observations and next observations $\mathcal{D}=\{(\boldsymbol{o}_{i},\boldsymbol{o}^{\prime}_{i})\}_{i=1}^{N}$ and only a small subset of it has action labels $\mathcal{D}_{a}=\{(\boldsymbol{o}_{i},\boldsymbol{o}^{\prime}_{i},\boldsymbol{a}_{i})\}_{i=1}^{N_{a}}$ , with $\lambda:=|\mathcal{D}_{a}|/|\mathcal{D}|\ll 1$. Note we drop the $i$ subscript when we do not need to refer to specific samples. In practice, $\boldsymbol{o}$ and $\boldsymbol{o}^{\prime}$ could be feature vectors extracted from the original observations (Cui et al., 2025), and $\boldsymbol{o}$ could be a stack of historical observations to account for partial observability (Bruce et al., 2024). The two datasets are assumed to come from the same distribution in most LAM work (Menapace et al., 2021; Bruce et al., 2024; Ye et al., 2024) [^1] Certain work explores mismatch case, e.g., $\mathcal{D}$ is human videos and $\mathcal{D}_{a}$ is robotics data (Chen et al., 2024b)..
Model. Recent LAM designs (Schmidt and Jiang, 2023; Bruce et al., 2024; Ye et al., 2024; Chen et al., 2024b; AgiBot-World, 2025) (see Figure 1) take a pair of consecutive observations $(\boldsymbol{o},\boldsymbol{o}^{\prime})$ as input, and output a latent $\boldsymbol{z}$ as well as a prediction of the next observation $\hat{\boldsymbol{o}}^{\prime}$. (We avoid calling $\boldsymbol{z}$ ‘latent action’ to avoid confusion with the real action.)
The LAM is typically decomposed into an inverse dynamics model (IDM) and forward dynamics model (FDM), implemented as deep neural networks, and trained via a reconstruction loss,
@@
  • The LAM latents can be used as input to a world model $\hat{T}$ trained to generate future frames, $\hat{T}(\boldsymbol{o}^{\prime}|\boldsymbol{o},\boldsymbol{z})$ (Sun et al., 2024; Chen et al., 2024b; Bruce et al., 2024; Menapace et al., 2021). These world models can generate higher-quality frames than the FDM in LAM, but have been used to only qualitatively interpret the meaning of the learned latents.
  • The LAM latents can be used as labels in the pre-training of a latent policy $\pi_{\text{latent}}(\hat{\boldsymbol{z}}|\boldsymbol{o})$ (similar to behavior cloning on actions). This policy may later be mapped to real actions, $\pi_{\text{map}}(\hat{\boldsymbol{a}}|\hat{\boldsymbol{z}})$  (Bruce et al., 2024; Schmidt and Jiang, 2023), or be followed by a fine-tuning phase on real actions (Schmidt and Jiang, 2023; Ye et al., 2024; Chen et al., 2024b).
In both use cases, it is hoped that the learned latents can be aligned with the true actions as closely as possible. As in Schmidt and Jiang (2023), “our hypothesis is … [this] may allow us to learn a latent action representation with a structure closely corresponding to the true action space”.
@@ 3.2 Linear LAM
Setting. We conduct analysis in the controlled Markov process (CMP) framework (Puterman, 2014) with $(\mathcal{O},\mathcal{A},T)$. At a given timestep, the agent receives observation $\boldsymbol{o}\in\mathcal{O}$ and takes action $\boldsymbol{a}\in\mathcal{A}$. The transition function describes the probability distribution over the next observation, $T(\boldsymbol{o}^{\prime}|\boldsymbol{o},\boldsymbol{a})$.
We consider vector observations, $\mathcal{O}=\mathbb{R}^{d_{o}}$ (which could be thought of as image observations processed with a pre-trained image encoder as in Cui et al. (2025); Chen et al. (2024b)). The action space $\mathcal{A}=\mathbb{R}^{d_{a}}$ has lower dimensionality than $\mathcal{O}$, i.e., $d_{a}\ll d_{o}$ [^2] Discrete actions can be represented as one-hot vectors within $\mathbb{R}^{d_{a}}$. While most of analysis (except that on bottleneck) should also apply to discrete actions, we confine our analysis in continuous actions.. These actions are mapped to create controllable changes in observation space via an action effect matrix $X\in\mathbb{R}^{d_{o}\times d_{a}}$, $\boldsymbol{q}=X\boldsymbol{a}$.
We assume linear LAM cannot access the action $\boldsymbol{a}$ during the training, but we will make use of it to evaluate the learned LAM in simulations. We choose the transition function to be an additive combination of controllable changes $\boldsymbol{q}\in\mathbb{R}^{d}$ (the state changes caused by the ego agent’s actions) and exogenous noise $\boldsymbol{\epsilon}\in\mathbb{R}^{d}$ (representing environmental stochasticity or other agent’s actions), i.e. with the next state generated via, $\boldsymbol{o}^{\prime}=\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon}$.
@@
Figure 2: Overview of linear LAM. Grey blocks represent learnable parameter matrices, giving rise to the predictive model $\hat{\boldsymbol{o}}^{\prime}=A\boldsymbol{o}+B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})$. The green parts illustrate linear LAM with data augmentation to reduce the amount of information the latent contains about the observation (in Section 4.3). The red parts illustrate linear LAM with auxiliary action prediction head to encourage the latent to focus on the controllable actions and suppress the noise signal (in Section 4.4).
@@
LAM $\text{LAM Objective}\coloneqq\max_{\boldsymbol{z}}\left[\mathcal{I}(\boldsymbol{z};\boldsymbol{a})-\mathcal{I}(\boldsymbol{z};\boldsymbol{\epsilon})-\mathcal{I}(\boldsymbol{z};\boldsymbol{o})\right],$ (5)
where $\mathcal{I}$ is mutual information. Hence, an ideal latent should contain all information about the action label, and no other parts of the environment, consistent with the argument from previous work – “biasing [the latent] towards simpler representations is likely preferrable” (Schmidt and Jiang, 2023).
Whilst mutual information is typically a challenging quantity to measure, in our linear model we can approximate this straightforwardly. After training the linear LAM via Eq. 4 and freezing the parameters, we fit three additional linear layers predicting $(\boldsymbol{q},\boldsymbol{\epsilon},\boldsymbol{o})$ from $\boldsymbol{z}$, to give $(\hat{\boldsymbol{q}},\hat{\boldsymbol{\epsilon}},\hat{\boldsymbol{o}})$ respectively. Note that, under the assumption that the variables $\mathbf{x}$ and $\mathbf{y}$ are independent multivariate Gaussian variables, the mutual information $\mathcal{I}(\mathbf{x},\mathbf{y})=-\dfrac{1}{2}\log\left(\frac{\lVert\hat{\mathbf{y}}-\mathbf{y}\rVert_{2}^{2}}{\mathbb{V}\text{ar}(\mathbf{y})}\right)$ where $\hat{\mathbf{y}}$ is the least squares estimate (LSE) of $\mathbf{y}$ based on $\mathbf{x}$, and $\mathbb{V}\text{ar}(\cdot)$ indicates the total variance (see e.g., Chapter 8 of (Johnson et al., 2002)) of a multivariate random variable. We define an evaluation metric that capture the objective of training linear LAM,
@@
Note that we get rid of the $\log(\cdot)$ function wrapped around the MSE loss to better calculate the values for this objective since the value range for the mutual information $[0,+\infty)$ is unbounded. This objective is maximized when $\boldsymbol{z}$ perfectly predicts $\boldsymbol{q}$ while containing no information about $\boldsymbol{\epsilon}$ and $\boldsymbol{o}$, resulting in the optimal value for LLO equal to $0+1+1=2$.
See the discussion on the gap between linear LAM and practical LAM in Appendix A.

### v3

@@ Abstract
Latent action models (LAMs) aim to learn action-relevant changes from unlabeled videos by compressing changes between frames as latents. However, differences between video frames can be caused by controllable changes as well as exogenous noise, leading to an important concern – do latents capture the changes caused by actions or irrelevant noise? This paper studies this issue analytically, presenting a linear model that encapsulates the essence of LAM learning, while being tractable. This provides several insights, including connections between LAM and principal component analysis (PCA), desiderata of the data-generating policy, and justification of strategies to encourage learning controllable changes using data augmentation, data cleaning, and auxiliary action-prediction. These findings are validated through numerical simulations, as well as experiments in more realistic settings. This investigation is the first to rigorously investigate how the structure of observations, actions, and noise influence LAM learning.
@@ 1 Introduction
Latent action models (LAMs) aim to infer controllable action changes from streams of image observations in an unsupervised manner (rybkin2018learning; menapace2021playable). This is valuable because action-labeled data is typically expensive to source, while unlabeled videos are abundant. Hence, it offers a route for embodied AI systems to learn from large unlabeled datasets, for instance using the inferred latent actions as targets for pre-training a policy, while a small amount of labeled data can be used to learn a mapping from latent to real action controls (ye2024lapa). This has proven effective in learning from videos of 2D video games, robotics, and even broadcast tennis footage (menapace2021playable; schmidt2023lapo; bruce2024genie; chen2024igor; ye2024lapa; sun2024deltadiffusion; cui2025dynamo; gao2025adaworld).
@@
The success of such LAM-based recipes relies on the inferred latent action labels mapping to the real control action signals of interest. However, there is concern that this may not always be the case. For example, there is an intuition that LAM’s inferred latent ‘actions’ simply compress differences between consecutive frames, even when control actions are not the cause of those differences (mccarthy2024towards). As such, LAMs may only succeed in domains where the cause of changes between observations can be fully attributed to the control action. A further point of dispute is whether a bottleneck is required (schmidt2023lapo).
@@
Figure 1: Linear LAM is an abstraction of the LAMs used in previous work. Inputting consecutive observation pairs $(\boldsymbol{o},\boldsymbol{o}^{\prime})$, the LAMs output the second observation via a reconstruction loss, $\lVert\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o}^{\prime}\rVert_{2}^{2}$. An information bottleneck tries to stop the direct copying of $\boldsymbol{o}^{\prime}$, with the expectation the latent $\boldsymbol{z}$ will correspond to the control action $\mathbf{a}$. Linear LAM captures the essence of LAM training whilst being analytically tractable. The diagrams of previous LAMs are copied from their original papers: LAPO (schmidt2023lapo), LAPA (ye2024lapa), Moto (chen2024moto), Genie (bruce2024genie), AdaWorld (gao2025adaworld), and Go-1 (agibot2025aigbot).
@@
To study these issues and more, this paper conducts a theoretical analysis of a linear version of LAM. By retaining the architecture of recent LAM models, but swapping deep neural network components with simpler linear layers, we preserve the fundamental challenge of LAM training in an analytically tractable form. Our analysis of linear LAM firstly provides precise insights into when inferred latent actions capture true control signals compared to noise, as well as what information is captured by different components within LAM. Surprisingly, our analysis also reveals additional issues not currently known to the LAM community – related to the over-parametrization property of LAM and the randomness of data-generation policy. Finally, we propose and study potential solutions to ameliorate these issues – data augmentation and predicting action as an auxiliary task.
@@
    Section 3 presents linear LAM, a tractable model preserving the essence of LAMs used in practice.
@@
    Section 4.1 shows linear LAM reduces to principal component analysis (PCA) on a mixture of controllable changes and exogenous noise, under certain assumptions. Our analysis justifies the practical use of LAM when the controllable action signals cause larger changes to observations than the exogenous noise.
@@
    Section 4.2 shows correlation between observations and actions decreases LAM’s focus on learning controllable changes. This suggests that higher randomness in data-generating policies benefits LAM’s learning.
@@
    Section 4.3 validates that performing data augmentation during LAM training can mitigate the over-parametrization issue and thus improve the semantics of the latent.
@@
    Section 4.4 finds that adding an action-prediction head encourages LAM to prioritize the learning of controllable changes for the latent.
  6. 6.
    Section 5 verifies that the main findings based on linear LAM still hold on more realistic LAMs.
@@ 2 Related Work
The study of learning representations of real actions in reinforcement learning (RL) has a long history. For instance, PG-RA (chandak2019learning) clusters actions based on the similarity of their impact on the state to improve generalization in the action space. LASER (allshire2021laser) learns latent actions through an encoder-decoder architecture trained to reconstruct real actions, resulting in higher learning efficiency in RL. TAP (jiang2022efficient) learns the latent action that can help to reconstruct the full trajectory (with state, action, and reward) condition on the state. EAR (hua2022simple) finds that the latent task embedding resulting from the training of multi-task policies turn out to be good action representations with a geometrically and semantically meaningful structure. AD3 (wang2024ad3) adopts an inverse dynamics model and a forward dynamics model to extract latent action, similar to popular LAMs, but conditions these models on the real actions.
@@
While the above papers learn action representations based on real actions, our paper focuses on learning latent actions without access to the real action labels. Removing the need for action labels during training is advantageous as it allows leveraging internet-scale video datasets in the pre-training stage (miech2019howto100m; chen2024panda; pei2025modeling; wang2023internvid) – for example unlabeled demonstrations of humans completing diverse tasks. While high-quality robotic datasets with action annotations exist (vuong2023open; fang2023rh20t; khazatsky2024droid; agibot2025aigbot), they remain limited in scale. Such datasets can be integrated into the semi-supervised learning framework to extract latent actions (nikulin2025distractor) or the policy fine-tuning stage (schmidt2023lapo; ye2024lapa). An alternative approach aims to extract pre-defined actions from observations using computer vision techniques (mendonca2023structured).
@@
Beginning from rybkin2018learning, LAMs have featured an information bottleneck or auto-encoder to allow learning in an unsupervised manner. ILPO (edwards2019imitating) learns latent actions along with the training of a policy that outputs the latent action and a world model that conditions on the latent action. menapace2021playable proposes a probabilistic action network that extracts a discrete action label and a continuous action variability embedding from consecutive observations. This network is trained jointly with an action decoder to generate video controlled by extracted actions. While these works adopt a bottleneck in the learning of latent actions, their training losses are complicated by involving policy learning or recurrent networks. LAPO (schmidt2023lapo) proposes a LAM design with an inverse dynamics model extracting a latent action and a forward dynamics model that reconstructs the next observation based on the latent action. This forms the template of the modern LAM. Many papers follow a similar architectures to LAPO with a discrete latent action space such as FICC (ye2022become), Genie (bruce2024genie), LAPA (ye2024lapa), Moto (chen2024moto), IGOR (chen2024igor), Go-1 (agibot2025aigbot), and GR001T N1 (nvidia2025gr00T). However, there is a debate whether discrete latent actions are better than continuous latent actions. For example, nikulin2025distractor finds that using continuous latent actions with a larger bottleneck can not only result in better predictability on real actions but also lead to better performance for downstream policies. We consider continuous latent actions in our linear LAM analysis while vector quantized latents are tested empirically in Section 5.
In contrast to the popularity of LAMs in recent foundation models for embodied AI, issues around the objective, learnability, and robustness of LAM have received limited attention. Consequently, our paper aims to investigate these issues.
@@ 3 Setup
This section broadly introduces the problem setting, goal and model used in recent LAM work in practice. We then more formally detail these for the linear model to be used in subsequent analysis. Finally, we outline the details of our simulation setup to be used in support of our later analysis.
@@ 3.1 LAMs in Practice
Setting. The setting tackled by recent LAM work assumes access to a large dataset of pairs of observations and next observations $\mathcal{D}=\{(\boldsymbol{o}_{i},\boldsymbol{o}^{\prime}_{i})\}_{i=1}^{N}$ and only a small subset of it has action labels $\mathcal{D}_{a}=\{(\boldsymbol{o}_{i},\boldsymbol{o}^{\prime}_{i},\boldsymbol{a}_{i})\}_{i=1}^{N_{a}}$ , with $\lambda:=|\mathcal{D}_{a}|/|\mathcal{D}|\ll 1$. Note we drop the $i$ subscript when we do not need to refer to specific samples. In practice, $\boldsymbol{o}$ and $\boldsymbol{o}^{\prime}$ could be feature vectors extracted from the original observations (cui2025dynamo), and $\boldsymbol{o}$ could be a stack of historical observations to account for partial observability (bruce2024genie). The two datasets are assumed to come from the same distribution in most LAM work (menapace2021playable; bruce2024genie; ye2024lapa) [^1] Certain work explores mismatch case, e.g., $\mathcal{D}$ is human videos and $\mathcal{D}_{a}$ is robotics data (chen2024igor)..
Model. Recent LAM designs (schmidt2023lapo; bruce2024genie; ye2024lapa; chen2024igor; agibot2025aigbot) (see Figure 1) take a pair of consecutive observations $(\boldsymbol{o},\boldsymbol{o}^{\prime})$ as input, and output a latent $\boldsymbol{z}$ as well as a prediction of the next observation $\hat{\boldsymbol{o}}^{\prime}$. (We subsequently avoid calling $\boldsymbol{z}$ ‘latent action’ to avoid confusion with the real action.)
LAMs are typically decomposed into an inverse dynamics model (IDM) and forward dynamics model (FDM), implemented as deep neural networks, and trained via a reconstruction loss,
@@
  • The LAM latents can be used as input to a world model $\hat{T}$ trained to generate future frames, $\hat{T}(\boldsymbol{o}^{\prime}|\boldsymbol{o},\boldsymbol{z})$ (sun2024deltadiffusion; chen2024igor; bruce2024genie; menapace2021playable). These world models can generate higher-quality frames than the FDM in LAM, but have been used to only qualitatively interpret the meaning of the learned latents.
[…]
In both use cases, it is hoped that the learned latents can be aligned with the true actions as closely as possible. As in schmidt2023lapo, “our hypothesis is … [this] may allow us to learn a latent action representation with a structure closely corresponding to the true action space”.
@@ 3.2 Linear LAM
Setting. We conduct analysis in the controlled Markov process (CMP) framework (puterman2014markov) with $(\mathcal{O},\mathcal{A},T)$. At a given timestep, the agent receives observation $\boldsymbol{o}\in\mathcal{O}$ and takes action $\boldsymbol{a}\in\mathcal{A}$. The transition function describes the probability distribution over the next observation, $T(\boldsymbol{o}^{\prime}|\boldsymbol{o},\boldsymbol{a})$.
We consider vector observations, $\mathcal{O}=\mathbb{R}^{d_{o}}$ (which could be thought of as image observations processed with a pre-trained image encoder as in cui2025dynamo; chen2024igor). The action space $\mathcal{A}=\mathbb{R}^{d_{a}}$ has lower dimensionality than $\mathcal{O}$, i.e., $d_{a}\ll d_{o}$ [^2] Discrete actions can be represented as one-hot vectors within $\mathbb{R}^{d_{a}}$. While most of analysis (except that on bottleneck) should also apply to discrete actions, we confine our analysis in continuous actions.. These actions are mapped to create controllable changes in observation space via an action effect matrix $X\in\mathbb{R}^{d_{o}\times d_{a}}$, $\boldsymbol{q}=X\boldsymbol{a}$.
We generally assume linear LAM cannot access the action $\boldsymbol{a}$ during the training, but we will make use of it to evaluate the learned LAM in simulations. We choose the transition function to be an additive combination of controllable changes $\boldsymbol{q}\in\mathbb{R}^{d}$ (the state changes caused by the ego agent’s actions) and exogenous noise $\boldsymbol{\epsilon}\in\mathbb{R}^{d}$ (representing environmental stochasticity or other agent’s actions), i.e. with the next state generated via, $\boldsymbol{o}^{\prime}=\boldsymbol{o}+\boldsymbol{q}+\boldsymbol{\epsilon}$.
@@
Appendix A summarizes the gap between linear and practical LAM.
Figure 2: Overview of linear LAM. Grey blocks represent learnable parameter matrices, giving rise to the predictive model $\hat{\boldsymbol{o}}^{\prime}=A\boldsymbol{o}+B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})$. Green illustrates linear LAM with data augmentation to reduce the amount of information the latent contains about the observation (in Section 4.3). Pink illustrate linear LAM with auxiliary action prediction to encourage the latent to focus on the controllable actions and suppress the noise signal (in Section 4.4).
@@
$\text{LAM Objective}\coloneqq\max_{\boldsymbol{z}}\left[\mathcal{I}(\boldsymbol{z};\boldsymbol{a})-\mathcal{I}(\boldsymbol{z};\boldsymbol{\epsilon})-\mathcal{I}(\boldsymbol{z};\boldsymbol{o})\right],$ (5)
where $\mathcal{I}$ is mutual information. Hence, an ideal latent should contain all information about the action label, and no other parts of the environment, consistent with intuitions of previous work – “biasing [the latent] towards simpler representations is likely preferrable” (schmidt2023lapo).
@@
$\text{Linear LAM Objective (LLO)}\coloneqq\max_{\boldsymbol{z}}\left[-\frac{\lVert\hat{\boldsymbol{q}}-\boldsymbol{q}\rVert_{2}^{2}}{\mathbb{V}\text{ar}(\boldsymbol{q})}+\frac{\lVert\hat{\boldsymbol{\epsilon}}-\boldsymbol{\epsilon}\rVert_{2}^{2}}{\mathbb{V}\text{ar}(\boldsymbol{\epsilon})}+\frac{\lVert\hat{\boldsymbol{o}}-\boldsymbol{o}\rVert_{2}^{2}}{\mathbb{V}\text{ar}(\boldsymbol{o})}\right].$ (6)
Note that we get rid of the $\log(\cdot)$ function wrapped around the MSE loss to better calculate the values for this objective since the value range for the mutual information $0,+\infty)$ is unbounded. This objective is maximized when $\boldsymbol{z}$ perfectly predicts $\boldsymbol{q}$ while containing no information about $\boldsymbol{\epsilon}$ and $\boldsymbol{o}$, resulting in the optimal value for LLO equal to $0+1+1=2$.